OpenStudy (anonymous):

An analytical procedure requires a solution of chloride ions. how many grams of Cacl2 must be dissolved to make 2.15 L of 0.0520 M Cl-? [Atomic mass: Ca=40.1 g/mol; Cl=35.5 g/mol]

OpenStudy (abb0t):

First, multiply 2.15 by the concentration, M, since M has units \(\sf \color{red}{\frac{mol}{L}}\). You can cancel out units: \(\sf \color{blue}{\cancel L \times \frac{mol}{\cancel L} = mol}\) now, multiply by the molar mass to get grams. And you see, that it's a 1:2 ratio of Ca:Cl which should tell you the grams.